A car is proceeding down a 1 in 10 incline at 65kph when it’s brakes are abruptly applied. How far does it skid along the road before stopping is the wheels lock and the coefficient of friction between the tires and the road is 0.7?


1
Expert's answer
2022-02-21T12:06:16-0500

Given:

"v=65\:\rm kph=18.1\: m/s"

"\sin\theta=1/10"

"\mu=0.7"


The acceleration of a car

"a=g(\sin\theta-\mu\cos\theta)\\
=9.8(1/10-0.7*\sqrt{9}/10)=-0.6\:\rm m/s^2"

The distance traveled by the car

"d=\frac{v_f^2-v_i^2}{2a}=\frac{0-18.1^2}{2*(-0.6)}=275\:\rm m"